The length of a rectangle is 2 inches more than its width. The area of the rectangle is 15 inches. What are the length and width of the rectangle?
Answer by mananth(16946)

Let the width be x
the length will be x+2
Area = L*W
=x(x+2)
15=x^2+2x
x^2+2x-15=0
x^2+5x-3x-15=0
x(x+5)-3(x+5)=0
(x+5)(x-3)=0
x=-5 or 3
width = 3 and length = 5